# Efficient Methods for Non-stationary Online Learning

### v2

@@ E.2.2 Proof of Theorem 7
Since we employ Adapt-ML-Prod and standard geometric covering schedule to ensemble the base-learners, then one can expect that $\mathtt{meta}\mbox{-}\mathtt{regret}\leq\mathcal{O}\left(\sqrt{k\cdot\lvert I\rvert\cdot\log T}\right)$. By Lemma 12, the base-regret is of order $\mathtt{base}\mbox{-}\mathtt{regret}\leq\mathcal{O}(\sqrt{k\cdot\lvert I\rvert})$. Combining these two bounds together, we finish the proof. ∎
@@ F.1 OGD and Dynamic Regret
[…]
@@
$\sum_{t=1}^{T}f_{t}(\mathbf{x}_{t})-\sum_{t=1}^{T}f_{t}(\mathbf{u}_{t})\leq\frac{7D^{2}}{4\eta}+\frac{D}{\eta}\sum_{t=2}^{T}\lVert\mathbf{u}_{t-1}-\mathbf{u}_{2}\rVert_{2}+\eta\sum_{t=1}^{T}\lVert\nabla f_{t}(\mathbf{x}_{t})\rVert_{2}^{2}$
@@
[…]
@@ F.2 Self-confident Tuning
[…]
$\mathbf{x}_{t}=\operatorname*{arg\,min}_{\mathbf{x}\in\mathcal{X}}\bigg{\{}\psi_{t}(\mathbf{x})+\sum_{\tau=1}^{t-1}f_{t}(\mathbf{x})\bigg{\}}$
@@
$\sum_{t=1}^{T}f_{t}(\mathbf{x}_{t})-f_{t}(\mathbf{u})\leq\frac{\psi(\mathbf{u})-\min_{\mathbf{x}\in\mathcal{X}}\psi(\mathbf{x})}{\eta_{T+1}}+\frac{1}{2\alpha}\sum_{t=1}^{T}\eta_{t}\lVert\nabla f_{t}(\mathbf{x}_{t})\rVert_{*}^{2}.$
@@
where the first inequality chooses $\mathbf{u}$ as the one-hot vector with all entries being 00 except the $i$-th one as $1$, and second inequality is by Lemma 23. ∎
@@ F.3 Facts on Geometric Covers
Let $[s_{p},s_{q}]\subseteq[T]$ be an arbitrary interval that starts from a marker $s_{p}$ and ends at another marker $s_{q}$. Then, we can find a sequence of consecutive intervals $I_{1}=[s_{i_{1}},s_{i_{2}}-1]$, $I_{2}=[s_{i_{2}},s_{i_{3}}-1]$, $\ldots$, $I_{v}=[s_{i_{v}},s_{i_{v+1}}-1]\in\widetilde{\mathcal{C}}$ such that $i_{1}=p$, $i_{v}\leq q<i_{v+1}$, and $v\leq\lceil\log_{2}(q-p+2)\rceil$.

### v3

@@ E.2.2 Proof of Theorem 7
[…]
@@ F.1 OGD and Dynamic Regret
Suppose that the regularizer $\psi:\mathcal{X}\mapsto\mathbb{R}$ is 1-strongly convex with respect to the norm $\|\cdot\|$. The dynamic regret of Optimistic Mirror Descent (OMD) whose update rule specified in (69) is bounded as follows:
@@
$\sum_{t=1}^{T}f_{t}(\mathbf{x}_{t})-\sum_{t=1}^{T}f_{t}(\mathbf{u}_{t})\leq\frac{7D^{2}}{4\eta}+\frac{D}{\eta}\sum_{t=2}^{T}\lVert\mathbf{u}_{t-1}-\mathbf{u}_{t}\rVert_{2}+\eta\sum_{t=1}^{T}\lVert\nabla f_{t}(\mathbf{x}_{t})\rVert_{2}^{2}$
@@
Suppose the regularizer $\psi:\mathcal{X}\mapsto\mathbb{R}$ is $1$-strongly convex with respect to the norm $\lVert\cdot\rVert$. The subsequent decisions $\mathbf{x}_{t+1},\mathbf{x}_{t}$ specialized in the OMD update rule (69) satisfy $\lVert\mathbf{x}_{t+1}-\mathbf{x}_{t}\rVert\leq\eta_{t}\lVert\nabla f_{t}(\mathbf{x}_{t})\rVert_{*}$.
@@ F.2 Self-confident Tuning
Suppose that the regularizer function $\psi:\mathcal{X}\mapsto\mathbb{R}$ is $\alpha$-strongly convex with respect to the norm $\|\cdot\|$. Let $f_{t}$ be a sequence of convex loss functions and $\psi_{t}(\mathbf{x})=\frac{1}{\eta_{t}}(\psi(\mathbf{x})-\min_{\mathbf{x}^{\prime}\in\mathcal{X}}\psi(\mathbf{x}^{\prime}))$, where $\eta_{t+1}\leq\eta_{t}$ holds for $t\in[T]$. Then the decision sequence $\mathbf{x}_{t}$ generated by the following FTRL update rule
$\mathbf{x}_{t}=\operatorname*{arg\,min}_{\mathbf{x}\in\mathcal{X}}\bigg{\{}\psi_{t}(\mathbf{x})+\sum_{\tau=1}^{t-1}f_{\tau}(\mathbf{x})\bigg{\}}$
@@
$\sum_{t=1}^{T}f_{t}(\mathbf{x}_{t})-\sum_{t=1}^{T}f_{t}(\mathbf{u})\leq\frac{\psi(\mathbf{u})-\min_{\mathbf{x}\in\mathcal{X}}\psi(\mathbf{x})}{\eta_{T+1}}+\frac{1}{2\alpha}\sum_{t=1}^{T}\eta_{t}\lVert\nabla f_{t}(\mathbf{x}_{t})\rVert_{*}^{2}.$
@@
where the first inequality chooses $\mathbf{u}$ as the one-hot vector with all entries being 0 except the $i$-th one as $1$, and second inequality is by Lemma 23. ∎
@@ F.3 Facts on Geometric Covers
Let $s_{p},s_{q}]\subseteq[T]$ be an arbitrary interval that starts from a marker $s_{p}$ and ends at another marker $s_{q}$. Then, we can find a sequence of consecutive intervals $I_{1}=[s_{i_{1}},s_{i_{2}}-1]$, $I_{2}=[s_{i_{2}},s_{i_{3}}-1]$, $\ldots$, $I_{v}=[s_{i_{v}},s_{i_{v+1}}-1]\in\widetilde{\mathcal{C}}$ such that $i_{1}=p$, $i_{v}\leq q<i_{v+1}$, and $v\leq\lceil\log_{2}(q-p+2)\rceil$.